two years back a father was 3 times as old as his son.two years hence, twice the fathers age will equal to 5 times that of his son's age.what are their present ages?
Answer by ankor@dixie-net.com(22740)

let f = father's present age
let s = son's present age
:
Write an equation and simplify it, for each statement
:
"two years back a father was 3 times as old as his son."
f - 2 = 3(s - 2)
f - 2 = 3s - 6
f = 3s - 6 + 2
f = 3s - 4
:
"two years hence, twice the fathers age will equal to 5 times that of his son's age."
2(f + 2) = 5(s + 2)
2f + 4 = 5s + 10
2f = 5s + 10 - 4
2f = 5s + 6
:
From the 1st equation we know f = (3s-4), replace f in the above equation
2(3s - 4) = 5s + 6
6s - 8 = 5s + 6
6s - 5s = 6 + 8
s = 14 yrs is son's present age
then
f = 3(14) - 4
f = 42 - 4
f = 38 yrs is father's present age
;
:
See if this works in the statement:
"two years hence, twice the fathers age will equal to 5 times that of his son's age."
2(38+2) = 5(14+2)
2(40) = 5(16); confirms our solutions of f=38 and s=14